# Over-the-air Federated Policy Gradient

### 31 Jan 2024 v2

@@ II-C Over-the-air Aggregation
In this paper, we consider the transmission from the agents to the central controller adopting the over-the-air aggregation, i.e., each agent broadcasts an analog signal carrying the information $M\hat{\nabla}J_{i}(\boldsymbol{\theta}^{k})$ to a common wireless channel. With transmitter synchronization and phase compensation [26], the central controller at time step $k$ can receive the following signal
@@ III Main Results
Under Assumption 1 and Assumption 2, if $\sigma_{h}^{2}\leq\frac{M(N+1)}{M-1}m_{h}^{2}$ and the stepsize $\alpha$ satisfies $\alpha\leq\frac{1}{m_{h}L}$, then
@@
Let $\epsilon>0$. If $\sigma_{h}^{2}\leq\frac{M(N+1)}{M-1}m_{h}^{2}$, the stepsize $\alpha$ satisfies $\alpha\leq\frac{1}{m_{h}L}$, and the epoch size $K$ and the batch size $M$ are chosen to satisfy $K=\mathcal{O}\left(\frac{1}{\epsilon}\right)$, $N=\mathcal{O}\left(\frac{1}{\sqrt{\epsilon}}\right)$ and $M=\mathcal{O}\left(\frac{1}{N\epsilon}\right)$, then the generated $\{\boldsymbol{\theta}^{k}\}$ satisfy $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]\leq\epsilon$. That is to say, by increasing $K$, $M$, and $N$, the proposed algorithm can converge to an $\epsilon$-approximate stationary point $\boldsymbol{\theta}$ satisfying that $\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta})\right\|^{2}\right]$ is arbitrarily small.
@@
The condition $\sigma_{h}^{2}\leq\frac{M(N+1)}{M-1}m_{h}^{2}$ puts a restriction on the channel statistics. Note that the more agents there are, the more easily the assumption can be satisfied. The convergence without this condition is shown in the following theorem, i.e., Theorem 2, which reveals the effect induced by the random channel states.
@@
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]\leq&\frac{2N(J\left(\boldsymbol{\theta}^{0})-J(\boldsymbol{\theta^{*}})\right)}{\alpha K(N+1)m_{h}}\\ &+\frac{\sigma_{h}^{2}V^{2}}{(N+1)m_{h}^{2}}+\frac{\sigma^{2}}{N(N+1)}.\end{split}$ (11)
@@ IV Simulation Results
We first consider that the channel gain $h_{i,k}$ follows Rayleigh distribution with $m_{h}=\sqrt{\frac{\pi}{2}}$ and $\sigma_{h}^{2}=\frac{4-\pi}{2}$. It can be easily verified that $\sigma_{h}^{2}\leq\frac{M(N+1)}{M-1}m_{h}^{2}$ holds for all $N$ and $M$. Fig. 3 and Fig. 3 present the averaged empirical cumulative reward and estimate of $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]$, respectively, with standard deviation. It can be seen that increasing both the batch size $M$ and the number of agents $N$ can improve the convergence performance. Specifically, Fig. 3 verifies Theorem 1 and shows that the proposed algorithm achieves linear speedup w.r.t. the number of agents $N$.
@@ Appendix C Proof of Theorem 1
If $\sigma_{h}^{2}\leq\frac{M(N+1)}{M-1}m_{h}^{2}$, we have
@@ Appendix D Proof of Theorem 2
$\begin{split}&\frac{\alpha(N+1)m_{h}}{2N}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]\\ \leq&\left(J(\boldsymbol{\theta}^{0})-J(\boldsymbol{\theta}^{*})\right)+\frac{\alpha\sigma_{h}^{2}KV^{2}}{2Nm_{h}}+\frac{\alpha m_{h}K\sigma^{2}}{2N^{2}}.\end{split}$ (23)
@@
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]\leq&\frac{2N\left(J(\boldsymbol{\theta}^{0})-J(\boldsymbol{\theta}^{*})\right)}{\alpha K(N+1)m_{h}}\\ &+\frac{\sigma_{h}^{2}V^{2}}{(N+1)m_{h}^{2}}+\frac{\sigma^{2}}{N(N+1)}.\end{split}$ (24)

### 26 Feb 2024 v3

@@ II-C Over-the-air Aggregation
In this paper, we consider the transmission from the agents to the central controller adopting the over-the-air aggregation, i.e., each agent broadcasts an analog signal carrying the information ^$\hat{\nabla}J_{i}(\boldsymbol{\theta}^{k})$ to a common wireless channel. With transmitter synchronization and phase compensation [26], the central controller at time step $k$ can receive the following signal
@@ III Main Results
Under Assumption 1 and Assumption 2, if $\sigma_{h}^{2}\leq(N+1)m_{h}^{2}$ and the stepsize $\alpha$ satisfies $\alpha\leq\frac{1}{m_{h}L}$, then
@@
Let $\epsilon>0$. If $\sigma_{h}^{2}\leq(N+1)m_{h}^{2}$, the stepsize $\alpha$ satisfies $\alpha\leq\frac{1}{m_{h}L}$, and the epoch size $K$ and the batch size $M$ are chosen to satisfy $K=\mathcal{O}\left(\frac{1}{\epsilon}\right)$, $N=\mathcal{O}\left(\frac{1}{\sqrt{\epsilon}}\right)$ and $M=\mathcal{O}\left(\frac{1}{N\epsilon}\right)$, then the generated $\{\boldsymbol{\theta}^{k}\}$ satisfy $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]\leq\epsilon$. That is to say, by increasing $K$, $M$, and $N$, the proposed algorithm can converge to an $\epsilon$-approximate stationary point $\boldsymbol{\theta}$ satisfying that $\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta})\right\|^{2}\right]$ is arbitrarily small.
@@
The condition $\sigma_{h}^{2}\leq(N+1)m_{h}^{2}$ puts a restriction on the channel statistics. Note that the more agents there are, the more easily the assumption can be satisfied. The convergence without this condition is shown in the following theorem, i.e., Theorem 2, which reveals the effect induced by the random channel states.
@@
$\begin{split}&\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]\\ \leq&\frac{2MNm_{h}\left(J(\boldsymbol{\theta}^{0})-J(\boldsymbol{\theta}^{*})\right)}{\alpha K(M(N+1)m_{h}^{2}+\sigma_{h}^{2})}+\frac{M\sigma_{h}^{2}V^{2}}{M(N+1)m_{h}^{2}+\sigma_{h}^{2}}\\ &+\frac{\sigma_{h}^{2}V^{2}}{M(N+1)m_{h}^{2}+\sigma_{h}^{2}}+\frac{Mm_{h}^{2}\sigma^{2}}{N(M(N+1)m_{h}^{2}+\sigma_{h}^{2})}.\end{split}$ (11)
@@ IV Simulation Results
We first consider that the channel gain $h_{i,k}$ follows Rayleigh distribution with $m_{h}=\sqrt{\frac{\pi}{2}}$ and $\sigma_{h}^{2}=\frac{4-\pi}{2}$. It can be easily verified that $\sigma_{h}^{2}\leq(N+1)m_{h}^{2}$ holds for all $N$ and $M$. Fig. 3 and Fig. 3 present the averaged empirical cumulative reward and estimate of $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]$, respectively, with standard deviation. It can be seen that increasing both the batch size $M$ and the number of agents $N$ can improve the convergence performance. Specifically, Fig. 3 verifies Theorem 1 and shows that the proposed algorithm achieves linear speedup w.r.t. the number of agents $N$.
@@ Appendix C Proof of Theorem 1
If $\sigma_{h}^{2}\leq(N+1)m_{h}^{2}<\frac{M(N+1)}{M-1}m_{h}^{2}$, we have
@@ Appendix D Proof of Theorem 2
$\begin{split}&\frac{\alpha(M(N+1)m_{h}^{2}+\sigma_{h}^{2})}{2MNm_{h}}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]\\ \leq&\left(J(\boldsymbol{\theta}^{0})-J(\boldsymbol{\theta}^{*})\right)+\frac{\alpha\sigma_{h}^{2}KV^{2}}{2Nm_{h}}+\frac{\alpha\sigma_{h}^{2}KV^{2}}{2MNm_{h}}+\frac{\alpha m_{h}\sigma^{2}K}{2N^{2}}.\end{split}$ (23)
@@
$\begin{split}&\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\|\nabla J(\boldsymbol{\theta}^{k})\right\|^{2}\right]\\ \leq&\frac{2MNm_{h}\left(J(\boldsymbol{\theta}^{0})-J(\boldsymbol{\theta}^{*})\right)}{\alpha K(M(N+1)m_{h}^{2}+\sigma_{h}^{2})}+\frac{M\sigma_{h}^{2}V^{2}}{M(N+1)m_{h}^{2}+\sigma_{h}^{2}}\\ &+\frac{\sigma_{h}^{2}V^{2}}{M(N+1)m_{h}^{2}+\sigma_{h}^{2}}+\frac{Mm_{h}^{2}\sigma^{2}}{N(M(N+1)m_{h}^{2}+\sigma_{h}^{2})}.\end{split}$ (24)